OpenStudy (anonymous):

if f(x) = x + 10 find (f of f^-1)(5)? is it all real numbers, 0, 5, or x?

OpenStudy (phi):

f^-1 means "f inverse" or the function that "undoes" the function f. example if f meant x^2 then f^-1(x) would be sqrt(x) so what happens if you do square(sqrt(4))? Let's see: sqrt(4)= 2. square(2)= 4 Does that give you an idea of what the answer is?

OpenStudy (anonymous):

so i need to find the sqr root of 10?

OpenStudy (phi):

I don't mean to confuse you. That was an example of a function and its inverse. The question is asking about a different function and its inverse. But the idea is that a function and its inverse together end up giving back the number they start with.

OpenStudy (anonymous):

haha im still lost? so its x?

OpenStudy (phi):

well, another way to answer this question is to find the inverse of f(x)=x+10 it is f^-1(x)= x-10 now form: f of f^-1 by replacing x in f(x)= x+10 with f^-1

OpenStudy (phi):

you get f of f^-1 = x and (f of f^-1)(5) = 5

OpenStudy (anonymous):

you dont understand.. im not smart at all math is my worst subject!

OpenStudy (phi):

behind all the mumbo jumbo are ideas that often are not too hard to understand. so think like this f(x) is a "thing" that you give an "x" to and it returns a "y" (x and y are numbers) f^-1(y) does the opposite. you give it y, and it returns x. so f(5) returns some number (call it y) f^-1(y) will return the 5

OpenStudy (phi):

Here is how it works with this problem f^-1(x)= x-10 and f^-1(5)= 5-10= -5 now put the -5 in f(x)=x+10 f(-5)= -5+10= 5 that means (f of f^-1)(5)= 5